OpenStudy (anonymous):

4(x - 2y + 2z) - (2x + 3y - 2z) for x = -1, y = 3, and z = 5

OpenStudy (anonymous):

4(x - 2y + 2z) - (2x + 3y - 2z) for x = -1, y = 3, and z = 5 can you please help me with this work i dont understand how to do this

OpenStudy (anonymous):

if x = -1 4 ( -1 -2y + 2z ) - ( -2 + 3y - 2z) -4 - 8y + 8z + 2 -3y + 2z -2 -11y + 10z

OpenStudy (anonymous):

Or if you use all of them in one, 4( -1 -( 2x3) + (2x5)) - ( -2 + 9 - 10) 4( -1-7+10) - ( -2 ) 8 + 2 10

OpenStudy (anonymous):

Forget about the first post i made, that one is wrong

OpenStudy (anonymous):

it is the wrong answer i still dont get how to do this work

OpenStudy (anonymous):

4(x - 2y + 2z) - (2x + 3y - 2z)=2x-11y+6z

OpenStudy (anonymous):

Now just put the values x = -1, y = 3, and z = 5 in the equation 2x-11y+6z

OpenStudy (anonymous):

-2-33+30=-5

OpenStudy (anonymous):

So the answer is -5

OpenStudy (anonymous):

\[2 (-1)-11\ 3+10\ 5 = 15 \]Simplify the expression \[4(x-2y+2z)-(2x+3y-2z)=2 x-11 y+10 z \] 2 x (-1) - 11 x (3) + 10 x 5 = 15